1.  A 10.0-liter flask contains  0.400 mole of H2, and 0.300 mole of  N2 at 20°C. 

(a) What is the pressure, in atmospheres, inside the flask?

(b) Calculate the mole fractions of the three gases?




1
Expert's answer
2021-10-15T02:58:39-0400

Moles of Nitrogen = 0.3


Moles of hydrogen = 0.4


Volume = 10 L


Temperature = 20 + 273 = 293 K


PV = nRT


Partial pressure of Nitrogen = 0.3×0.0821×208/10= 0.512atm


Partial pressure of hydrogen= 0.4×0.0821×208/10= 0.68atm


Total pressure = 0.512+0.68= 1.192atm